OpenStudy (anonymous):

f(x)=x/x+1 & g(x)= x/1-x which of the following statements is true?

OpenStudy (anonymous):



OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@zordoloom

OpenStudy (anonymous):

Have you tired solving any of the following?

OpenStudy (anonymous):

yes, i got B but I think im wrong

OpenStudy (anonymous):

Do you mind showing how you solved it?

OpenStudy (anonymous):

\[f(g(x))=f(\frac{x}{1-x})=\frac{\frac{x}{1-x}}{\frac{x}{1-x}+1}\] then a raft of algebra to get rid of the compound fraction multiply top and bottom by \(1-x\) is the quickest way to do it